OpenStudy (anonymous):

How many marbles, each with a volume of 36 cubic centimeters, are needed to fill in a cylindrical vase with a radius of 6 centimeters and a height of 28 centimeters?

OpenStudy (triciaal):

marble = sphere volume of a sphere = 4/3 *pi* r^3 = 36 volume of a cylinder = pi*r^2*h where r is the radius and h is the height

OpenStudy (anonymous):

how would i find the height? @triciaal

OpenStudy (triciaal):